OpenStudy (anonymous):

4. What happens to the force between two masses if the mass of one object is doubled?

OpenStudy (rajat97):

can you please draw a diagram

OpenStudy (anonymous):

\[F_1=G\frac{Mm_1}{r^2}\] if the distance between the masses stays the same, and \[m_2=2m_1\] \[F_2=G\frac{Mm_2}{r^2}=G\frac{M(2m_1)}{r^2}=2\Bigg(G\frac{Mm_1}{r^2}\Bigg)=2F_1\] Doubling the mass of one of the objects doubles the force between them.